# FIRE [Uncaptioned image]: Flexible Integration of Data Quality Ratings for Effective Pre-Training

### v1

@@ Abstract
Selecting high-quality data can significantly improve the pre-training efficiency of large language models (LLMs). Existing methods often rely on heuristic techniques and single quality signals, limiting their ability to comprehensively evaluate data quality. In this work, we propose FIRE, a flexible and scalable framework for integrating multiple data quality raters, which allows for a comprehensive assessment of data quality across various dimensions. FIRE aligns multiple quality signals into a unified space, and integrates diverse data quality raters to provide a comprehensive quality signal for each data point. Further, we introduce a progressive data selection scheme based on FIRE that iteratively refines the selection of high-quality data points, balancing computational complexity with the refinement of orthogonality. Experiments on the SlimPajama dataset reveal that FIRE consistently outperforms other selection methods and significantly enhances the pre-trained model across a wide range of downstream tasks, with a 2.9% average performance boost and reducing the FLOPs necessary to achieve a certain performance level by more than half.
FIRE [Uncaptioned image]: Flexible Integration of Data Quality Ratings
for Effective Pre-Training
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance by pre-training on trillions of tokens using large-scale Transformers. However, constrained by scaling laws Kaplan et al. (2020), LLMs are rapidly nearing their capacity and data limits. Consequently, the pursuit of improving performance has shifted towards optimizing the quality of pre-training data.
@@
Recent studies indicate that effective data selection can significantly enhance the convergence speed and generalization capability of pre-trained models Engstrom et al. (2024); Wettig et al. (2024). Existing methods predominantly rely on heuristic techniques, including rule-based filtering Rae et al. (2021); Raffel et al. (2020), deduplication Abbas et al. (2023); Tirumala et al. (2024), and assessing proximity to high-quality corpora Xie et al. (2023). Novel approaches have also emerged, enhancing the evaluation of pre-training data quality by querying authoritative LLMs whether the texts conform to specific criteria Wettig et al. (2024); Sachdeva et al. (2024).
@@
Figure 1: Few-shot downstream performance with respect to pre-training FLOPs for random, FIRE, and FIRE progressive.
@@
Intuitively, evaluating the quality of a text requires analyzing multiple dimensions. Nevertheless, the methods mentioned above assess data quality based on individual aspects, lacking a comprehensive evaluation of the data’s overall quality. The success of querying LLMs lead to the intuitive idea of combining various quality rating criteria into a single prompt to solicit comprehensive quality signals from authoritative LLMs. However, experimental findings show that this approach considerably weakens the evaluative performance of LLMs, as the excessive number of rules makes it challenging for LLMs to follow (further details can be found in Appendix E.2). The challenge of adhering to multiple rules highlights the necessity for a more sophisticated strategy to effectively integrate multiple quality signals.
@@
In this work, we propose FIRE, a Flexible and scalable framework for Integrating multiple data quality Raters, designed to enable Effective pre-training of LLMs. Initially, we propose an alignment method to address the issue of inconsistent ratings from multiple raters. The method involves the probability that the data in each quantile is of higher quality (win rate) compared to a reference subset as the aligned rating. By fitting a win-rate-quantile curve for each rater, we effectively map the ratings from multiple raters into a unified rating space. Next, to derive a comprehensive signal representing the overall data quality, we integrate the aligned ratings of multiple raters, considering both the intrinsic reliability and orthogonality of the raters. Further, we introduce a progressive data selection scheme based on FIRE that iteratively refines the selection of high-quality data points, balancing computational complexity with the refinement of orthogonality.
@@
Our pre-training experiments on the SlimPajama Soboleva et al. (2023) dataset demonstrate that by applying integrated ratings from multiple raters, our method achieves superior results across various downstream tasks. Figure 1 illustrates that FIRE substantially improves the pre-trained model, offering a 2.9% average increase in performance while simultaneously reducing the FLOPs required to attain a specific performance level by half. We summarize our main contributions as follows:
@@
(3) Our pretraining experiments on the SlimPajama dataset demonstrate that FIRE effectively boosts the performance of the pre-trained model by 2.9% on average, while also more than halving the FLOPs necessary to attain a particular performance level.
@@ 2.1 Overview of the method
We develop a method called FIRE that flexibly integrates multiple raters to effectively assess the comprehensive quality of data, which contains two processes: (a) Rating Alignment and (b) Rater Integration. The overall framework of FIRE is illustrated in Figure 2.
@@
We then integrate the aligned ratings of multiple raters, considering both the intrinsic reliability and orthogonality of the raters. We constructs an orthogonality graph and calculates centrality through PageRankPage et al. (1999) to quantify the independence among raters. The integrated rating is given by:
@@ 2.2 Rating Alignment
It’s critical to align ratings from multiple data quality raters to achieve a credible integrated rater. This involves standardizing ratings to a consistent scale and addressing the significant variances in raters’ high-quality thresholds (i.e., the score thresholds distinguishing data that positively contribute to pre-training). Appendix C.1 provides further analyses on the importance of rating alignment.
Given a pre-training dataset $\mathcal{D}_{t}$ and multiple raters $R_{1},R_{2},\ldots,R_{n}$, we propose a method to consolidate these ratings into a unified rating space:
@@
Uniformly sample a subset $\mathcal{D}_{r}$ from the pre-training data $\mathcal{D}_{t}$.
@@
Randomly sample a subset $\mathcal{D}_{ij}$ from each interval, ensuring that $|\mathcal{D}_{ij}|=|\mathcal{D}_{r}|$. Use GPT-4o Islam and Moushi (2024) to evaluate how the $\mathcal{D}_{ij}$ samples impact pre-training in comparison to the reference dataset $\mathcal{D}_{r}$. Then, calculate the win rate $w_{ij}$ for each interval $j$:
@@ 2.3 Rater Integration
0:  The entire dataset $\mathcal{D}_{t}$; Decay factor $\eta$; Initial number of parts $n$; Part multiplication factor $\beta$; Maximum number of parts $n_{\text{max}}$; Desired dataset size $k$
0:  $\mathcal{D}_{s}$: Selected data subset
1:  Calculate and sort integrated ratings for $\mathcal{D}_{t}$
2:  Reduce the data to $\eta\%$ of its original size using decay factor $\eta$
3:  while size of $\mathcal{D}_{t}$ $\geq k$ do
4:     Divide the data into $n$ parts based on the quantiles of the integrated ratings
5:     for each part $P_{i}$ do
6:        Calculate the overall orthogonality in $P_{i}$
7:        Derive refined integrated ratings $\mathcal{S}_{P_{i}}$
8:     end for
9:     Sort the data based on the new integrated ratings $\mathcal{S}_{P_{i}}$
10:     Reduce the data to $\eta\%$ of its original size
11:     if $n\times\beta\leq n_{\text{max}}$ then
12:        $n\leftarrow n\times\beta$
13:     else
14:        $n\leftarrow n_{\text{max}}$
15:     end if
16:  end while
17:  $\mathcal{D}_{s}$leftarrow top k elements from $\mathcal{D}_{t}$
@@
Baseline (Integration method) Comp. Rater 52.9 24.3 81.2 22.0 62.0 30.9 59.2 50.1 47.8
@@ 4.1 Experimental setup
We use SlimPajama Soboleva et al. (2023) as the selection pool, with a total scale of 627B. And we employ the Llama Touvron et al. (2023) tokenizer to divide the entire dataset into sequences of length 1024. During the data selection process, we select the top portion of data with the highest rating. For integrating different raters, we carry out experiments based on the four single raters of QuRating Wettig et al. (2024): Writing Style, Facts and Trivia, Educational Value, and Required Expertise. For progressive data selection (FIRE Progressive), we set $\eta=60,\beta=20$. For model training, we train a model with 1.3B parameters for 10,000 steps (equivalent to 20B tokens), with bfloat16 format during both training and testing.
@@
We utilize lm-evaluation-harness Gao et al. (2021) to assess our model’s performance across eight downstream tasks: ARC-E Clark et al. (2018), ARC-C Clark et al. (2018), SciQ Welbl et al. (2017), LogiQA Liu et al. (2020), BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), PIQA Bisk et al. (2020), and WinoGrande Sakaguchi et al. (2021). We employ in-context learning for the evaluation, selecting enough demonstrations to fill the window length of 1024 tokens for each task. Standard accuracy metrics are reported for all tasks.
@@
In addition to comparing our method with four single raters from Qurating, we also compare it with the following methods: (1) Random: randomly selecting data from the original training corpus with a uniform distribution. (2) DSIR Xie et al. (2023): utilizing importance sampling for data selection, and we chose the Wikipedia and Books datasets as target domains. (3) Density Sachdeva et al. (2024): using KDE to estimate data density in the training corpus and employing inverse sampling. (4) ASK-LLM Sachdeva et al. (2024): using a comprehensive prompt to label high quality data, and train a T5-based classifier.
@@ 4.2 Main Results
Comparing the FIRE results to those of the individual raters, our rating integration method demonstrates significant improvements, with an average score increase of up to 2.9% over the best single rater. Additionally, our method outperforms Random, QuRating and other data selection methods, validating the high quality of the FIRE selected data.
@@
We observe that as the number of integrated raters increases, the overall effect gradually improves. The average score of FIRE (3 Raters) is 49.6, surpassing the score of the integration of two raters. Furthermore, when four raters are integrated, the average score increases to 50.0. This indicates that our rater integration method is scalable: incorporating a broader range of raters not only provides a more comprehensive evaluation of the samples, but also allows for more precise calibration of the importance of each metric.
@@
When compared to other integration methods, FIRE proves to be more effective. As shown in Table 1, Comprehensive Rater using a merged multi-dimension prompt results in an average score that is even lower than the single-dimension rater Facts and Trivia. This suggests that GPT-4o still falls short in assessing data quality from a broad perspective. Both Mix Criteria and Max Criteria amplify a sample’s best-performing dimension, but their scores are still inferior to ours, indicating that a comprehensive evaluation is more beneficial. Average simply calculates the mean of all ratings; however, we believe that different dimensions should not equally impact model performance, and experimental results show an increase of 3.3% over Average.
@@ 4.3 Analysis
From Figure 3, we can find that: (1) Rating alignment is a crucial step. We note that without rating alignment, the score drops by 3.3%, even falling below the direct average. As we previously detailed, the alignment allows for better comparability between the ratings from different raters, making their integration more reasonable. (2) Both orthogonality and intrinsic reliability can further enhance model performance, with the impact of orthogonality being relatively significant (a drop of 2% w/o Orth.), while the improvement from Intrinsic Reliability is rather subtle (a drop of 0.2% w/o Rel.) (3) The combination of all components yields the best results. This implies that these methods of integrating the ratings are complementary. By superimposing both methods, we can achieve a rating that more accurately reflects the actual contribution of the sample to the pre-training.
@@
Figure 4: The in-context learning results with respect to pre-training FLOPS on four downstream tasks: ARC-E, ARC-C, SciQ, and HellaSwag. FLOPS are approximated as $6\times model\ size\times dataset\ size$ Kaplan et al. (2020).
Figures 1 and 6 show how the model’s performance on downstream tasks evolves with the pre-training FLOPS. In terms of average score, our method outperforms the random baseline by 2.9%. From the training efficiency perspective, our method reduces training FLOPS by 60%. In addition, our method shows a significant advantage in the ARC-E/C and SciQ tasks, consistently scoring higher than the random baseline. However, on the HellaSwag task, our method’s performance is similar to the random baseline and does not consistently surpass it. One possible explanation is that HellaSwag is an especially challenging dataset, which makes it difficult to discern performance differences on the 1.3B scale model.
@@
Moreover, we assess the focus of each rater in comparison to the FIRE. Based on the four dimensions in QuRating and a comprehensive evaluation dimension, we determine the percentage of high-quality data selected by each rater in each dimension. We pick 1K data points at random from the top 20B data that are selected by each method. Then we use GPT-4o to assess each dimension six times, taking the average as the final evaluation result. Refer to the Appendix D.3 for more details. As we can see from Table 3: (1) From a comprehensive perspective, FIRE shows the best results, as evidenced by our performance in downstream tasks. (2) In terms of each dimension, FIRE consistently achieves relatively high accuracy, demonstrating that the data selected by this method maintains high quality across all dimensions.
Based on the results of previous experiments, compared to single raters, our method could potentially select more diverse samples, which is beneficial for model training. In order to validate this hypothesis, we take the top-rated 1K texts from each method, and encode these texts using Sentence-T5 Ni et al. (2022). Subsequently, we reduce dimensions using t-SNE Van der Maaten and Hinton (2008), with the visualization results displayed in Figure 6(b). In the latent semantic space, the top data selected by single raters tends to be clustered, whereas data selected by FIRE spans a broader range, ensuring the diversity of the selected data. Moreover, as seen from previous case studies, the data selected by FIRE also maintains high quality, suggesting that FIRE has achieved a good balance between diversity and quality.

### v2

# FIRE [Uncaptioned image]: Flexible Integration of Data Quality Ratings for Effective Pretraining

@@ Abstract
Selecting high-quality data can significantly improve the pretraining efficiency of large language models (LLMs). Existing methods generally rely on heuristic techniques and single-quality signals, limiting their ability to evaluate data quality comprehensively. In this work, we propose FIRE, a flexible and scalable framework for integrating multiple data quality raters, which allows for a comprehensive assessment of data quality across various dimensions. FIRE aligns multiple quality signals into a unified space, and integrates diverse data quality raters to provide a comprehensive quality signal for each data point. Further, we introduce a progressive data selection scheme based on FIRE that iteratively refines the selection of high-quality data points. Experiments on the SlimPajama dataset reveal that FIRE outperforms other data selection methods and significantly enhances the pretrained model across a wide range of downstream tasks, with a 2.9% average performance improvement over Random and reducing the FLOPs necessary to achieve a certain performance level by more than half.
FIRE [Uncaptioned image]: Flexible Integration of Data Quality Ratings for Effective Pretraining
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance by utilizing large-scale Transformers to pretrain on trillions of tokens. However, due to the constraints imposed by scaling laws Kaplan et al. (2020), LLMs are quickly nearing their capacity and data limits. As a result, efforts to improve LLM performance have increasingly concentrated on optimizing the quality of pretraining data.
@@
Numerous studies indicate that effective data selection can significantly enhance the convergence speed and generalization capability of LLMs Engstrom et al. (2024); Wettig et al. (2024); Gao et al. (2025). Traditional methods predominantly rely on heuristic techniques, such as rule-based filtering Rae et al. (2021); Raffel et al. (2020), deduplication Abbas et al. (2023); Tirumala et al. (2024), and assessing proximity to high-quality corpora Xie et al. (2023). Additionally, some work has focused on improving the evaluation of pretraining data quality by querying authoritative LLMs to determine whether the texts meet specific criteria Wettig et al. (2024); Sachdeva et al. (2024).
@@
Figure 1: Few-shot downstream performance with respect to pretraining FLOPs for random, FIRE, and FIRE progressive.
@@
Intuitively, assessing the quality of a text involves analyzing it across multiple dimensions. Nevertheless, the methods mentioned above evaluate data quality based on individual aspects, lacking a comprehensive assessment of the data’s overall quality. Building on this limitation, the success of querying LLMs Sachdeva et al. (2024) has inspired the straightforward idea of merging various quality rating criteria into a single prompt to collect comprehensive quality signals from authoritative LLMs. However, experimental findings show that this approach considerably weakens the performance of LLMs, as the excessive number of rules makes it challenging for LLMs to follow (further details can be found in Appendix E.2). The challenge of adhering to multiple rules underscores the necessity for a more sophisticated strategy to integrate various quality signals effectively.
@@
In this paper, we propose FIRE, a Flexible and scalable framework for Integrating multiple data quality Raters, designed to enable Effective pretraining of LLMs. Initially, we introduce an alignment method to tackle the issue of inconsistent ratings from multiple raters. This method involves ranking the data based on the scores from the original raters and then partitioning it into quantiles. We assess the probability that the data within each quantile is of higher quality (win rate) compared to a reference subset, using this probability as the aligned rating. By fitting a win-rate-quantile curve for each rater, we effectively map the ratings from multiple raters into a unified rating space. Subsequently, to derive a comprehensive signal representing the overall data quality, we integrate the aligned ratings of multiple raters, considering both the intrinsic reliability and orthogonality of the raters. Further, we introduce a progressive data selection scheme based on FIRE that iteratively refines the selection of high-quality data points, balancing computational complexity with the refinement of orthogonality.
@@
Extensive experiments on the SlimPajama Soboleva et al. (2023) dataset demonstrate that by applying integrated ratings from multiple raters, our method achieves superior results across a variety of downstream tasks. Figure 1 illustrates that FIRE significantly enhances the pretrained model, providing an average performance increase of 2.9% while also reducing the FLOPs needed to achieve a specific performance level by more than half. We summarize our main contributions as follows:
@@
(3) Extensive experiments on the SlimPajama dataset demonstrate that FIRE enhances the pretrained model’s performance by an average of 2.9% while reducing the FLOPs required to achieve a certain performance level by more than half.
@@ 2.1 Overview of the Method
We propose FIRE, a method that flexibly integrates multiple raters to comprehensively evaluate data quality. It involves two key processes: (a) Rating Alignment and (b) Rater Integration. Figure 2 illustrates the overall framework of FIRE.
@@
We then integrate the aligned ratings of multiple raters, considering both the intrinsic reliability and orthogonality of the raters. We construct an orthogonality graph and calculate centrality through PageRankPage et al. (1999) to quantify the independence among raters. The integrated rating is given by:
@@ 2.2 Rating Alignment
Aligning ratings from multiple data quality raters is crucial for achieving a credible integrated rater. This process involves standardizing ratings to a consistent scale and eliminating the significant differences in raters’ high-quality thresholds, which are the score thresholds that distinguish data contributing positively to pretraining. Appendix C.1 offers further analysis on the importance of rating alignment.
Given a pretraining dataset $\mathcal{D}_{t}$ and multiple raters $R_{1},R_{2},\ldots,R_{n}$, we propose a method to consolidate these ratings into a unified rating space:
@@
Uniformly sample a subset $\mathcal{D}_{r}$ from the pretraining data $\mathcal{D}_{t}$.
@@
Randomly sample a subset $\mathcal{D}_{ij}$ from each interval, ensuring that $|\mathcal{D}_{ij}|=|\mathcal{D}_{r}|$. Use GPT-4o Islam and Moushi (2024) to evaluate how the $\mathcal{D}_{ij}$ samples impact pretraining in comparison to the reference dataset $\mathcal{D}_{r}$. Then, calculate the win rate $w_{ij}$ for each interval $j$:
@@ 2.3 Rater Integration
1:  Input: The entire dataset $\mathcal{D}_{t}$; Decay factor $\eta$; Initial number of parts $n$; Part multiplication factor $\beta$; Maximum number of parts $n_{\text{max}}$; Desired dataset size $k$
2:  Output: $\mathcal{D}_{s}$: Selected data subset
3:  Calculate and sort integrated ratings for $\mathcal{D}_{t}$
4:  Reduce the data to $\eta\%$ of its original size using decay factor $\eta$
5:  while size of $\mathcal{D}_{t}$ $\geq k$ do
6:     Divide the data into $n$ parts based on the quantiles of the integrated ratings
7:     for each part $P_{i}$ do
8:        Calculate the overall orthogonality in $P_{i}$
9:        Derive refined integrated ratings $\mathcal{S}_{P_{i}}$
10:     end for
11:     Sort the data based on the new integrated ratings $\mathcal{S}_{P_{i}}$
12:     Reduce the data to $\eta\%$ of its original size
13:     $n\leftarrow\min(n\times\beta,n_{\text{max}})$
14:  end while
15:  $\mathcal{D}_{s}$leftarrow top k elements from $\mathcal{D}_{t}$
@@
Baseline (Integration method, 4 Raters) Comp. Rater 52.9 24.3 81.2 22.0 62.0 30.9 59.2 50.1 47.8
@@ 4.1 Experimental setup
We use SlimPajama Soboleva et al. (2023) as the selection pool, with a total scale of 627B. And we employ the Llama Touvron et al. (2023) tokenizer to divide the entire dataset into sequences of length 1024. During the data selection process, we select the top portion of data with the highest rating. For integrating different raters, we carry out experiments based on the four single raters of QuRating Wettig et al. (2024): Writing Style, Facts and Trivia, Educational Value, and Required Expertise. For progressive data selection (FIRE Progressive), we set $\eta=60,\beta=20$. For model training, we train a model with 1.3B parameters for 10,000 steps (equivalent to 20B tokens) and a 3B model for 200B tokens, with bfloat16 format during both training and testing.
@@
We utilize lm-evaluation-harness Gao et al. (2021) to assess the models’ performance across eight downstream tasks: ARC-E Clark et al. (2018), ARC-C Clark et al. (2018), SciQ Welbl et al. (2017), LogiQA Liu et al. (2020), BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), PIQA Bisk et al. (2020), and WinoGrande Sakaguchi et al. (2021). We employ in-context learning for the evaluation, selecting enough examples to fill the window length of 1024 tokens for each task. Standard accuracy metrics are reported for all tasks.
@@
In addition to comparing FIRE with four single raters from Qurating, we also compare it with the following methods: (1) Random: randomly selecting data from the original training corpus. (2) DSIR Xie et al. (2023): utilizing importance sampling for data selection, and we chose Wikipedia and Books as target domains. (3) Density Sachdeva et al. (2024): using KDE to estimate data density in the training corpus and employing inverse sampling. (4) ASK-LLM Sachdeva et al. (2024): using a comprehensive prompt to label high-quality data, and train a T5-based classifier.
@@ 4.2 Main Results
Comparing the FIRE results to those of the individual raters, our rating integration method demonstrates significant improvements, with an average score increase of up to 1.9% over the best single rater. Additionally, our method outperforms Random, QuRating and other data selection methods, validating the high quality of the data selected by FIRE.
@@
We observe that as the number of integrated raters increases, the overall effect gradually improves. The average score of FIRE with three raters surpasses that of the integration of two raters, and further increases when four raters are integrated. This indicates that our rater integration method is scalable: incorporating a broader range of raters not only provides a more comprehensive evaluation of the samples but also allows for a better understanding of the importance of each metric.
@@
When compared to other integration methods, FIRE proves to be more effective. As shown in Table 1, Comprehensive Rater with a multi-dimension prompt results in an average score that is even lower than the single-dimension rater Facts and Trivia. This suggests that GPT-4o still falls short in assessing data quality from a broad perspective. Both Mix Criteria and Max Criteria are inferior to FIRE, indicating that a comprehensive evaluation is more beneficial. Average simply calculates the mean of all ratings; however, different dimensions should not have an equal impact on model performance, and the experimental results of FIRE demonstrate an improvement over Average.
@@ 4.3 Analysis
From Figure 3, we can find that: (1) Rating alignment is a crucial step. We note that without rating alignment, the score drops by 3.3%, even falling below the direct average. As we previously detailed, the alignment allows for better comparability between the ratings from different raters, making their integration more reasonable. (2) Both orthogonality and intrinsic reliability can further enhance model performance, with the impact of orthogonality being relatively significant (a drop of 2% w/o Orth.), while the improvement from Intrinsic Reliability is rather subtle (a drop of 0.2% w/o Rel.) (3) The combination of all components yields the best results. This implies that these methods of integrating the ratings are complementary. By superimposing both methods, we can achieve a rating that more accurately reflects the actual contribution of the sample to the pretraining.
@@
Figure 4: The in-context learning results with respect to pretraining FLOPS on four downstream tasks: ARC-E, ARC-C, SciQ, and HellaSwag. FLOPS are approximated as 6 × model size × dataset size Kaplan et al. (2020)
Figures 1 and 4 show how the model’s performance on downstream tasks evolves with the pretraining FLOPS. In terms of average score, our method outperforms the random baseline by 2.9%. From the training efficiency perspective, our method reduces training FLOPS to achieve a certain performance level by more than half. In addition, our method shows a significant advantage in the ARC-E/C and SciQ tasks, consistently scoring higher than the random baseline. However, on the HellaSwag task, our method’s performance is similar to the random baseline and does not consistently surpass it. One possible explanation is that HellaSwag is an especially challenging dataset, which makes it difficult to discern performance differences on the 1.3B scale model.
@@
Moreover, we analyze the data properties that FIRE focuses on compared to other raters. Based on the four dimensions in QuRating and a comprehensive evaluation dimension, we determine the percentage of high-quality data selected by each rater in each dimension. We pick 1000 data points at random from the top 20B data that are selected by each method. Then we use GPT-4o to assess each dimension six times, taking the average as the final evaluation result. Refer to the Appendix D.3 for more details. As we can see from Table 3: (1) From a comprehensive perspective, FIRE shows the best results, as evidenced by our performance in downstream tasks. (2) In terms of each dimension, FIRE consistently achieves relatively high accuracy, demonstrating that the data selected by this method maintains high quality across all dimensions.
Based on the results of previous experiments, compared to single raters, our method could potentially select more diverse samples, which is beneficial for model training. In order to validate this hypothesis, we take the top-rated 1000 texts from each method and encode these texts using Sentence-T5 Ni et al. (2022). Subsequently, we reduce dimensions using t-SNE Van der Maaten and Hinton (2008), with the visualization results displayed in Figure 6(b). In the latent semantic space, the top data selected by single raters tends to be clustered, whereas data selected by FIRE spans a broader range, ensuring the diversity of the selected data. Moreover, as seen from previous case studies, the data selected by FIRE also maintains high quality, suggesting that FIRE has achieved a good balance between diversity and quality.